# A Cognitive Paradigm Approach to Probe the Perception-Reasoning Interface in VLMs

### v3

@@ 7.2 Impact of Description Quality on Reasoning
The results, presented in Table 4, were revealing. Providing high-quality descriptions dramatically improved the reasoning accuracy of VLMs that struggled when using their own descriptions. Llama-Vision-11B, for example, improved from 53.4% (Table 2) to 84.17%, and Llama-Vision-90B from 55.1% to 90.98%. This provides strong evidence that the reasoning capabilities of these models are significantly underestimated by end-to-end evaluations; their primary limitation lies in generating accurate perceptual representations. Further illustrating this sensitivity to description source quality, Table A.4 in the Appendix details a comparison using components generated by Pixtral-12B versus GPT-4o.
Remarkably, this approach also enabled text-only LLMs to perform the visual reasoning task effectively. Models such as Phi-4 (14B) achieved 91.98% accuracy, surpassing the average human baseline, while several Qwen models also exceeded 90%. This demonstrates that: (1) High-quality textual descriptions can serve as effective surrogates for visual input, enabling modality transfer for reasoning tasks. (2) The Componential Analysis paradigm, particularly when coupled with controlled descriptive input, serves as a powerful tool for isolating and evaluating the core symbolic reasoning abilities of both VLMs and LLMs, independent of their integrated perceptual systems. These findings strongly reinforce the conclusion that improving visual perception is paramount for enhancing end-to-end visual reasoning in many current models.
Model Accuracy
Llava:7b 80.56
Llava:34b 81.56
Llama3.2-vision:11b 84.17
Llama3.2-vision:90b 90.98
Deepseek-r1:14b 87.98
Gemma2:27b 88.98
Qwen2.5:7b 90.38
Phi4:14b 91.98
Qwen2.5:32b 92.79
Qwen2.5:14b 92.99
7.3 Analysis per Commonsense Category
@@ 8 Discussion
This research employed a cognitively-inspired lens to investigate the mechanisms underlying visual reasoning in VLMs, using the challenging Bongard Problem paradigm. By structuring the evaluation around paradigms reflecting human cognitive strategies – holistic processing (DVRL), deductive reasoning (DRL), and analytical decomposition (CA) – we aimed to move beyond accuracy metrics towards understanding how these models process visual information and derive abstract rules.
@@
Our findings highlight the critical interdependence of perception and reasoning. While human cognition often involves rapid perceptual grouping (‘gist’ perception Biederman (1987); Li et al. (2002)) followed by more focused attention and deliberation Posner (1980); Kahneman (2011), our results suggest many current VLMs struggle significantly with the initial perceptual stage. The strong performance of advanced models like GPT-4o and Gemini 2.0 under the Componential Analysis paradigm – where visual complexity is reduced to textual descriptions – indicates powerful downstream reasoning capabilities. However, the dramatic performance drop for many open-source models in this same paradigm (when using their own descriptions) points squarely to a perception bottleneck.
@@
The ablation studies reinforce this conclusion. Providing high-quality external rules (Section 7.1) isolates rule application fidelity, while providing high-quality external descriptions (Section 7.2) isolates reasoning from perception. The results clearly show that reasoning performance is significantly enhanced when the perceptual input is reliable. This suggests that for many systems, failures attributed to “reasoning” may actually stem from noisy or inaccurate initial visual representations. From a cognitive perspective, this is akin to reasoning failures caused by misperception rather than flawed logic.
@@
The progressive performance improvement observed across paradigms (DVRL < DRL < CA for top models on natural images) suggests that structuring the reasoning process, particularly by separating distinct cognitive stages like rule extraction and application, or perception and reasoning, benefits current VLM architectures. This resonates with findings on Chain-of-Thought prompting Wei et al. (2022) but our staged paradigms offer a different way to structure and diagnose the process, particularly relevant for multi-modal tasks involving complex visual comparisons.
@@
Furthermore, the Componential Analysis paradigm demonstrates a viable approach for evaluating multi-image reasoning tasks even with single-image input models. By converting images to text, it allows models without inherent multi-image capacity to engage with tasks like BPs. It also successfully bridged the modality gap, enabling LLMs to perform sophisticated visual reasoning when provided with rich descriptions, achieving high performance. This highlights the potential of symbolic representations derived from vision as a powerful interface for general reasoning engines.
However, the poor performance on the synthetic Bongard Logo dataset serves as a crucial counterpoint. It suggests limitations either in the models’ ability to generalize beyond natural image statistics present in their pre-training data, or in the effectiveness of textual descriptions (generated by current VLMs) for capturing the essential abstract features of synthetic geometric patterns. This requires further investigation into how VLMs represent and reason about different types of visual structures.
@@ 10 Conclusion
Understanding the cognitive mechanisms enabling visual reasoning in AI is paramount for building truly intelligent systems. This paper introduced a novel evaluation framework, grounded in principles of human cognition, designed to dissect the processes VLMs employ when solving challenging visual reasoning tasks like Bongard Problems. Through three distinct paradigms – Direct Visual Rule Learning, Deductive Rule Learning, and Componential Analysis – we systematically probed the interplay between perception and reasoning.
Our key findings reveal both the impressive potential and critical limitations of current VLMs. Advanced models demonstrate strong reasoning capabilities, achieving high accuracy, particularly when the task is structured or when reasoning over high-quality textual representations derived from images. However, our analysis, particularly through Componential Analysis and targeted ablations, identifies a significant perception bottleneck in many contemporary open-source models, where failures in visual processing hinder downstream reasoning. Our framework uniquely enables the evaluation of multi-image reasoning tasks even for models architecturally constrained to single images, and successfully bridges the modality gap to evaluate text-only LLMs, further highlighting the importance of the perception-to-representation pipeline.
The diagnostic insights provided by our cognitively-inspired framework underscore the need for focused research on enhancing the fidelity and robustness of visual perception modules within VLMs. By offering tools to disentangle perception from reasoning and to evaluate performance across different cognitive strategies, this work contributes to a deeper understanding of VLM capabilities and limitations, guiding future efforts towards developing AI with more comprehensive and human-like visual intelligence.
@@ Appendix A.2 Attention and Memory in Visual Reasoning
While our study primarily focuses on the interplay between perception and reasoning, the roles of attention and memory are also implicitly present in our paradigms. The DVRL paradigm likely engages VLM “visual attention” mechanisms Bahdanau et al. (2016) to identify salient features across the image set, akin to human holistic processing Biederman (1987); Li et al. (2002). DRL relies on the model’s ability to “memorize” the extracted rule, involving processes related to working memory Baddeley (2012) and internal representation storage Squire (1992). Although not directly measured, their involvement is inherent. Future work could explore these aspects more explicitly, perhaps via attention map analysis Vaswani et al. (2017) or probing memory representations Vaishnav and Serre (2023).
@@ A.3.1 Bongard OpenWorld Dataset
We utilize a subset of 500 test cases from the Bongard OpenWorld dataset Wu et al. (2024). The full dataset contains 1001 samples, each with 7 positive and 7 negative real-world images distinguished by a “commonsense” rule. Our evaluation set was created by taking the first 250 samples and generating two test cases from each (one positive query, one negative query), resulting in 500 balanced test cases. Specific sample IDs used will be released.
A.3.2 Commonsense Value Categories
@@
Table A.1: Commonsense ID Categories and Examples in Bongard OpenWorld dataset Wu et al. (2024).
A.3.3 Commonsense Value Distribution in Our Subset
@@
Table A.2: Distribution of Commonsense ID Categories in the 500 Bongard-OpenWorld test cases used.
@@ A.3.4 Dataset Availability
Bongard OpenWorld: https://rujiewu.github.io/Bongard-OpenWorld.github.io/. Bongard-HOI: Check original paper Jiang et al. (2022). Bongard Logo: Check original paper Nie et al. (2020).
@@ A.4.1 Model Details
VLMs: GPT-4o; Gemini 2.0; Pixtral-12B; Llama-Vision-3.2 (11B, 90B); LLaVA (Llama-2 based; 7B, 13B, 34B); LLaVA-Llama3-8B. Text-Only LLMs (for Ablation 7.2): Phi-4 (14B) Abdin et al. (2024); Qwen2.5 (7B, 14B, 32B) Yang et al. (2024); Deepseek-r1 (32B, 70B) Guo et al. (2025); Gemma2 (27B) Team et al. (2024).
@@ A.4.3 Evaluation Metrics
  • Semantic Similarity: Cosine similarity of OpenAI embeddings (‘text-embedding-3-large‘) between descriptions/rules. Inspired by Risch et al. (2021).
@@ A.6.2 Performance on Bongard-HOI
A.6.3 Performance on Bongard Logo
(Results ( 50% accuracy) discussed in main text Section 6.2)
@@ A.6.5 Componential Analysis Results by Commonsense Category
ID Concept Category GPT 4o Gemini 2.0
@@
Table A.5: Overall accuracy of GPT 4o and Gemini 2.0 on the Bongard OpenWorld test set using Componential Analysis, broken down by Commonsense ID.
@@ A.6.7 Detailed Error Analysis Examples
Table A.7: Error Analysis: Examples of Bongard-OpenWorld cases misclassified by both GPT-4o and Gemini 2.0 in Componential Analysis. Captions indicate the ground truth rule Wu et al. (2024). Reasoning based on analyzing GPT-4o’s generated analysis, rule, and query description.

### v4

@@ 7.2 Impact of Description Quality on Reasoning
Model Accuracy
Llava:7b [40] 80.56
Llava:34b [40] 81.56
Llama3.2-vision:11b [39] 84.17
Llama3.2-vision:90b [39] 90.98
Deepseek-r1:14b [44] 87.98
Gemma2:27b [45] 88.98
Qwen2.5:7b [43] 90.38
Phi4:14b [42] 91.98
Qwen2.5:32b [43] 92.79
Qwen2.5:14b [43] 92.99
@@
The results, presented in Table 5, were revealing. Providing high-quality descriptions dramatically improved the reasoning accuracy of VLMs that struggled when using their own descriptions. Llama-Vision-11B, for example, improved from 53.4% (Table 1) to 84.17%, and Llama-Vision-90B from 55.1% to 90.98%. This provides strong evidence that the reasoning capabilities of these models are significantly underestimated by end-to-end evaluations; their primary limitation lies in generating accurate perceptual representations. Further illustrating this sensitivity to description source quality, Table A.5 in the Appendix details a comparison using components generated by Pixtral-12B.
Remarkably, this approach also enabled text-only LLMs to perform the visual reasoning task effectively. Models such as Phi-4 (14B) achieved 91.98% accuracy, surpassing the average human baseline, while several Qwen models also exceeded 90%. This demonstrates that: (1) High-quality textual descriptions can serve as effective surrogates for visual input, enabling modality transfer for reasoning tasks. (2) The CA paradigm, particularly when coupled with controlled descriptive input, serves as a powerful tool for isolating and evaluating the core symbolic reasoning abilities of both VLMs and LLMs, independent of their integrated perceptual systems. These findings strongly reinforce the conclusion that improving visual perception is paramount for enhancing end-to-end visual reasoning in many current models.
@@ 8 Discussion
This research leveraged a cognitively-inspired framework to dissect the visual reasoning mechanisms of VLMs. By evaluating performance across paradigms mirroring human strategies (holistic, deductive, analytical), we moved beyond accuracy to probe how models process visuo-conceptual information. A central finding is the critical role, and frequent limitation, of visual perception. While advanced VLMs possess powerful reasoning capabilities, unlocked effectively by the Componential Analysis (CA) paradigm when using high-fidelity textual descriptions, many models are bottlenecked by their inability to reliably extract and represent relevant visual information.
@@
Our findings highlight the critical interdependence of perception and reasoning. While human cognition often involves rapid perceptual grouping (‘gist’ perception [11, 49]) followed by more focused attention and deliberation [50, 51], our results suggest many current VLMs struggle significantly with the initial perceptual stage. The strong performance of advanced models like GPT-4o and Gemini 2.0 under the CA paradigm – where visual complexity is reduced to textual descriptions – indicates powerful downstream reasoning capabilities. However, the dramatic performance drop for many open-source models in this same paradigm (when using their own descriptions) points squarely to a perception bottleneck.
The success of the CA paradigm, achieving SOTA on diverse benchmarks like Bongard-OpenWorld, Bongard-HOI, and notably Winoground, is particularly insightful. Its strength appears to lie in decoupling perception from reasoning via task-agnostic description generation. Unlike end-to-end models or context-dependent CoT approaches that might conflate perceptual errors with reasoning failures or overfit to linguistic cues, CA first aims to build a comprehensive, independent representation of each visual input in text. This rich textual world model then allows powerful LLMs (whether multi-modal or text-only) to apply their sophisticated reasoning abilities effectively, leading to robust and generalizable performance across different reasoning types (abstraction in BPs, compositionality in Winoground). This suggests modularity, where specialized perception modules generate rich symbolic representations for general-purpose reasoning engines, might be a highly effective architecture.
@@
The ablation studies reinforce this conclusion. Providing external rules (Section 7.1) isolates rule application fidelity, while providing high-quality external descriptions (Section 7.2) isolates reasoning from perception. The results clearly show that reasoning performance is significantly enhanced when the perceptual input is reliable. This suggests that for many systems, failures attributed to “reasoning” may actually stem from noisy or inaccurate initial visual representations. From a cognitive perspective, this is akin to reasoning failures caused by misperception rather than flawed logic.
@@
The progressive performance improvement observed across paradigms (DVRL < DRL < CA for top models on natural images) suggests that structuring the reasoning process, particularly by separating distinct cognitive stages like rule extraction and application, or perception and reasoning, benefits current VLM architectures. This resonates with findings on Chain-of-Thought prompting [29] but our staged paradigms offer a different way to structure and diagnose the process, particularly relevant for multi-modal tasks involving complex visual comparisons.
@@
Furthermore, the CA paradigm demonstrates a viable approach for evaluating multi-image reasoning tasks even with single-image input models. By converting images to text, it allows models without inherent multi-image capacity to engage with tasks like BPs. It also successfully bridged the modality gap, enabling LLMs to perform sophisticated visual reasoning when provided with rich descriptions, achieving high performance. This highlights the potential of symbolic representations derived from vision as a powerful interface for general reasoning engines.
@@ 10 Conclusion
This paper introduced a cognitively-inspired evaluation framework to dissect the perception-reasoning interface in VLMs using diverse visual reasoning tasks. Through three paradigms (DVRL, DRL, CA), we systematically analyzed VLM processing. Our key contribution is the Componential Analysis (CA) approach, which decouples perception into task-agnostic textual descriptions, enabling robust reasoning by powerful language models. This method achieves new state-of-the-art results on challenging benchmarks requiring different reasoning styles: Bongard-OpenWorld (abstraction), Bongard-HOI (interaction), and Winoground (compositionality). This success highlights the power of rich symbolic representations derived from vision. Concurrently, our framework diagnoses a critical perception bottleneck limiting many current VLMs, whose reasoning abilities are significantly unlocked when provided with high-fidelity descriptions (confirmed via ablations and LLM evaluations). Our work provides valuable diagnostic tools and suggests that enhancing perceptual fidelity and exploring decoupled perception-reasoning architectures are crucial steps towards achieving general, robust visual intelligence in AI.
@@ Appendix A.2 Attention and Memory in Visual Reasoning
While our study primarily focuses on the interplay between perception and reasoning, the roles of attention and memory are also implicitly present in our paradigms. The DVRL paradigm likely engages VLM “visual attention” mechanisms [52] to identify salient features across the image set, akin to human holistic processing [11, 49]. DRL relies on the model’s ability to “memorize” the extracted rule, involving processes related to working memory [53] and internal representation storage [54]. Although not directly measured, their involvement is inherent. Future work could explore these aspects more explicitly, perhaps via attention map analysis [55] or probing memory representations [56].
@@ A.3.1 Bongard OpenWorld Dataset
We utilize a subset of 500 test cases from the Bongard OpenWorld dataset [8]. The full dataset contains 1001 samples, each with 7 positive and 7 negative real-world images distinguished by a “commonsense” rule. Our evaluation set was created by taking the first 250 samples and generating two test cases from each (one positive query, one negative query), resulting in 500 balanced test cases. Specific sample IDs used will be released.
A.3.1.1 Commonsense Value Categories
@@
Table A.1: Commonsense ID Categories and Examples in Bongard OpenWorld dataset [8].
A.3.1.2 Commonsense Value Distribution in Our Subset
@@
Table A.2: Distribution of Commonsense ID Categories in the 500 Bongard-OpenWorld test cases used in our evaluation subset.
A.3.2 Bongard-HOI Dataset
To assess generalizability on natural images with a different reasoning focus (human-object interactions), we used the Bongard-HOI dataset [9]. We evaluated performance on its four standard test splits, defined by object/action novelty:
  • sosa: seen object, seen action
  • soua: seen object, unseen action
  • uosa: unseen object, seen action
  • uoua: unseen object, unseen action
The original splits vary significantly in size and balance (e.g., sosa: 200 pos/200 neg queries; soua: 2236 pos/1348 neg; uosa: 660 pos/660 neg; uoua: 695 pos/695 neg). For consistent cross-split evaluation in this work, we created balanced subsets by sampling 100 test cases from each of the four splits, ensuring an equal distribution of 50 positive and 50 negative query images per split. This resulted in a total evaluation set of 400 samples for Bongard-HOI (100 per split), used for the results reported in Table 2.
A.3.3 Winoground Dataset
To test performance on fine-grained visio-linguistic compositional reasoning, we utilized the Winoground dataset [10]. This dataset comprises 400 samples specifically designed to challenge compositional understanding. Each sample contains a pair of minimally contrastive images ($I_{0},I_{1}$) and a corresponding pair of minimally contrastive captions ($C_{0},C_{1}$), requiring models to correctly match image $I_{0}$ to caption $C_{0}$ and image $I_{1}$ to caption $C_{1}$. We used all 400 samples provided in the standard dataset release for our Winoground evaluations reported in Section 6.3 and Table 3.
@@ A.3.4 Dataset Availability
Details on the specific subsets and samples used in our evaluations will be released upon publication.
@@ A.4.1 Model Details
VLMs: GPT-4o; Gemini 2.0; Pixtral-12B; Llama-Vision-3.2 (11B, 90B); LLaVA (Llama-2 based; 7B, 13B, 34B); LLaVA-Llama3-8B. Text-Only LLMs (for Ablation 7.2): Phi-4 (14B) [42]; Qwen2.5 (7B, 14B, 32B) [43]; Deepseek-r1 (32B, 70B) [44]; Gemma2 (27B) [45].
@@ A.4.3 Evaluation Metrics
  • Semantic Similarity: Cosine similarity of OpenAI embeddings (‘text-embedding-3-large‘) between descriptions/rules. Inspired by [57].
A.4.4 Winoground Score Calculation using Componential Analysis
This section defines the calculation of Winoground [10] scores (text_score, image_score, group_score) within our Componential Analysis (CA) paradigm (Section 6.2).
In the standard Winoground task, a sample $i$ consists of two images $I_{0,i},I_{1,i}$ and two captions $C_{0,i},C_{1,i}$, where $(C_{0,i},I_{0,i})$ and $(C_{1,i},I_{1,i})$ are the ground truth correct pairs. Models are typically evaluated based on a scoring function $s(C,I)$ indicating the match between a caption and an image.
In our CA paradigm, the reasoning model (Stage 2) does not access images $I_{0,i},I_{1,i}$ directly. Instead, it operates on textual image descriptions $D_{0,i},D_{1,i}$ generated in Stage 1. The model is prompted to make explicit choices about the best match between descriptions and captions. Let $Choice_{C}(D_{k},\{C_{0},C_{1}\})$ denote the caption ($C_{0}$ or $C_{1}$) chosen by the model as the best match for description $D_{k}$. Similarly, let $Choice_{D}(C_{k},\{D_{0},D_{1}\})$ denote the description ($D_{0}$ or $D_{1}$) chosen for caption $C_{k}$.
The scores for each sample $i$ in the dataset $W$ (where $N=|W|=400$) are calculated as follows:
1. Text Score ($f_{CA}$): This measures if the correct caption is selected for each image description. We use an indicator function $\mathbb{I}[\cdot]$ which is 1 if the condition inside is true, and 0 otherwise.
$f_{CA}(i)=\mathbb{I}\Big{[}\begin{gathered}Choice_{C}(D_{0,i},\{C_{0,i},C_{1,i}\})=C_{0,i}\\ and\\ Choice_{C}(D_{1,i},\{C_{0,i},C_{1,i}\})=C_{1,i}\end{gathered}\Big{]}$ (1)
This score is 1 only if the model correctly identifies the caption for both description $D_{0,i}$ and description $D_{1,i}$.
2. Image Score ($g_{CA}$): This measures if the correct image description is selected for each caption.
$g_{CA}(i)=\mathbb{I}\Big{[}\begin{gathered}Choice_{D}(C_{0,i},\{D_{0,i},D_{1,i}\})=D_{0,i}\\ and\\ Choice_{D}(C_{1,i},\{D_{0,i},D_{1,i}\})=D_{1,i}\end{gathered}\Big{]}$ (2)
This score is 1 only if the model correctly identifies the description for both caption $C_{0,i}$ and caption $C_{1,i}$.
3. Group Score ($h_{CA}$): This requires all associations within the sample to be correct.
$h_{CA}(i)=f_{CA}(i)\land g_{CA}(i)$ (3)
Equivalently, $h_{CA}(i)=1$ if and only if $f_{CA}(i)=1$ and $g_{CA}(i)=1$.
@@ A.6.2 Performance on Bongard-HOI
A.6.3 Winoground Performance Context
Model / Strategy Text Image Group
Gemini (Baseline) 30.75 26.00 25.00
Gemini + DDCoT 45.00 25.00 23.75
Gemini + CCoT 22.50 33.00 20.75
Gemini + CoCoT 40.00 32.50 27.75
Gemini 2.0 + CA (Ours) 71.91 48.71 42.01
Table A.4: Performance comparison on Winoground (400 samples). CA refers to our Componential Analysis paradigm. Other results use Gemini Pro Vision with different prompting strategies.
To contextualize the performance of our Componential Analysis (CA) paradigm applied to Gemini 2.0 on Winoground (reported in Section 6.2), we also ran evaluations using Gemini Pro Vision with several prompting strategies. Table A.4 shows these comparative results on the 400-sample Winoground set used. While advanced CoT methods like DDCoT and CoCoT improve over the baseline for Gemini Pro Vision, the CA paradigm applied to Gemini 2.0 achieves competitive scores, particularly on the text metric, demonstrating its effectiveness.
@@ A.6.5 Componential Analysis Results by Commonsense Category
ID Concept Category GPT-4o (%) Gemini 2.0 (%)
@@
Table A.6: Overall accuracy (%) of GPT-4o and Gemini 2.0 on the Bongard-OpenWorld test set using Componential Analysis, broken down by Commonsense ID category. Performance variations highlight differing model strengths on specific concept types.
@@ A.6.7 Detailed Error Analysis Examples
Table A.8: Error Analysis: Examples of Bongard-OpenWorld cases misclassified by both GPT-4o and Gemini 2.0 in Componential Analysis. Captions indicate the ground truth rule [8]. Reasoning based on analyzing GPT-4o’s generated analysis, rule, and query description.